OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

OpenStudy (anonymous):

y = mx +C m = (y-C)/x

OpenStudy (anonymous):

m is slope

OpenStudy (anonymous):

move all terms to one side and the slope = (- coefficient of x/coefficient of y) = -5/-2=5/2

OpenStudy (anonymous):

Im sorry, im not understanding.

OpenStudy (anonymous):

The slope of a line y = mx + c is m so in your case -2y = -10-5x, so : 2y =5x+10 so y = (5/2)x + 5 so m = 5/2 so slope is 5/2

OpenStudy (anonymous):

Thanks, I understand now :)

OpenStudy (anonymous):

you're welcome :) vote me up

OpenStudy (anonymous):

@moyle Please avoid doing all the work and give out the final answer as a reply to the question. Thanks! look at the link http://openstudy.com/code-of-conduct